OpenStudy (anonymous):

Can anyone explain a simple method for performing partial fraction decomposition relative to Laplace Transformations?

OpenStudy (anonymous):

why yes i can good sir!! but it works best with an example

OpenStudy (anonymous):

gimme a sec. i shall return.

OpenStudy (anonymous):

no particular example actually, i'm just trying to figure out a quick method for breaking down annoying laplace transforms :P

OpenStudy (anonymous):

lol, i hear ya, i hate those things. annoy my balls off.

OpenStudy (anonymous):

\[(s+3) \div s ^{2}(s+1)(s+2)\]

OpenStudy (anonymous):

that would be a nice one to see performed

OpenStudy (anonymous):

done and done, sec. responding to another Q

OpenStudy (anonymous):

kk, no rush

OpenStudy (anonymous):

ok, so we have s+3/(s^2(s+1)(s+2)) yes? so we can set up our equation as follows: s+3/(s^2(s+1)(s+2))= A/(s)+ Bx+C/s^2 + D/(s+1) + E/(s+2)

OpenStudy (anonymous):

sec

OpenStudy (zarkon):

Just use WolframAlpha http://alturl.com/idnwj ;)

OpenStudy (anonymous):

ok, so remember that any time you have an s^2 in the denominator, you need to have As+B in the numerator, instead of just a single variable. and if it's a s^3 in the denominator it's an As^2+Bs+C in the numerator, etc, etc.

OpenStudy (anonymous):

i noticed you used A/s + (Bs+C)/s^2 is that necessary? seems redundant

OpenStudy (anonymous):

yeah, it is necessary, otherwise you miss a term in your differential equation. also any time you have something like (s+2)^2 you have to have A/(s+2) + Bs+C/(s+2)^2

OpenStudy (zarkon):

oh..no!!

OpenStudy (zarkon):

A/s+B/s^2 is fine

OpenStudy (anonymous):

but (Bs+C)/s^2 = B/s + C/s^2 then A/s + B/s combine. when you transform back from laplace, you factor out the constant anyways. These are the things that confuse me lol

OpenStudy (anonymous):

is it?! ahh!!! i suck tonight.lol. i apologize for misleading

OpenStudy (anonymous):

zarkon wins all the internets for catching that.

OpenStudy (anonymous):

I may just have to look for an app for my ti-89. such a thing should exist

OpenStudy (zarkon):

when you have something like \[(s+2)^3\] you get \[\frac{A}{s+1}+\frac{B}{(s+2)^2}+\frac{C}{(s+2)^3}\]

OpenStudy (zarkon):

use the expand command

OpenStudy (anonymous):

I never could get expand to work the way I needed it too. Time to fish out that booklet now where the hell is it at lol. anyways, thanks for the help